OpenStudy (anonymous):

find derivative of \[f(x)=\frac{ 4 }{ \sqrt{x} }-\frac{ 3 }{ x } + \frac{ 9 }{ x^4 }\]

OpenStudy (chaise):

Remembering that: \[\frac{1}{a^{b}}=a^{-b}\] \[f(x)=4x^{-\frac{1}{2}}-3x ^{-1}+9x ^{-4}\] You can now use the power rule. If you are still having trouble I might be able to help.

OpenStudy (anonymous):

so would it be? \[-\frac{ 3 }{ 2x^2 }+\frac{ 2 }{ 3x }-\frac{ 5 }{ 36x }\]

OpenStudy (anonymous):

actually that can't be the answer in another homework I have multiple choice answers which are A.\[f'(x)=\frac{ 2 }{ x ^{3/2} }+\frac{ 3 }{ x^2 } \frac{ 36 }{ x^5 }\] B. \[f'(x)= -\frac{ 2 }{ x ^{3/2} } -\frac{ 3 }{ x^2 } - \frac{ 36 }{ x^3 }\] C.\[f'(x)=-2\sqrt{x}+\frac{ 3 }{ x^2 }-\frac{ 36 }{ x^3 }\] D.\[f'(x)=\frac{ 2 }{ x ^{1/2} }-\frac{ 3 }{ x^2 }-\frac{ 36 }{ x^5 }\]

OpenStudy (anonymous):

ah i got it \[f'(x)=-\frac{ 2 }{ x ^{3/2} }+\frac{ 3 }{ x^2 }-\frac{ 36 }{ x^5 }\]